# zip2zip: Inference-Time Adaptive Vocabularies for Language Models via Token Compression

### v1

@@ Abstract
Tokenization efficiency plays a critical role in the performance and cost of large language models (LLMs), yet most models rely on static tokenizers optimized on general-purpose corpora. These tokenizers’ fixed vocabularies often fail to adapt to domain- or language-specific inputs, leading to longer token sequences and higher computational costs. We introduce zip2zip, a framework that enables LLMs to dynamically adjust the token vocabulary at inference time, allowing for fewer generated tokens and thus faster inference. zip2zip consists of three key components: (1) a tokenizer based on Lempel-Ziv-Welch (LZW) compression that incrementally merges co-occurring tokens into reusable hypertokens on the fly; (2) an embedding layer that computes embeddings for newly formed hypertokens at runtime; and (3) a causal language modeling variant that trains the model to operate on hypertokenized, compressed sequences. We show that an existing LLM can be zip2zip-fied in 10 GPU-hours via parameter-efficient finetuning. The resulting zip2zip LLMs effectively learn to use hypertokens at inference time, reducing input and output sequence length by 20–60%, with significant improvements in inference latency. Code will be released at epfl-dlab/zip2zip.
@@ 1 Introduction
Large language models (LLMs) have shown impressive versatility across a broad spectrum of tasks and domains (Brown et al., 2020; Bubeck et al., 2023), including biomedical tests (Nori et al., 2023), mathematical reasoning (Frieder et al., 2023), programming (Jiang et al., 2024), and multiple human languages. A critical underlying component of this flexibility is the tokenizer, which defines the model’s vocabulary and governs how raw text is converted into token sequence fed to the model. The efficiency of the tokenization scheme—i.e., how compactly a text is represented as tokens—has significant impact on model performance. In particular, a more compact tokenization yields three key benefits: (1) larger effective context window; (2) lower computational (and thus monetary) cost; and (3) shorter response times.
Despite its importance, the tokenizer used in most LLMs produces a fixed, static vocabulary using algorithms such as Byte Pair Encoding (Sennrich et al., 2016) over large-scale, general-purpose web corpora. While this globally optimized vocabulary performs reasonably well on average, it often fails to adapt to domain-specific or language-specific distributions (Ahia et al., 2023; Petrov et al., 2023), where the text distribution diverges significantly from the pretraining data. The resulting mismatch leads to longer token sequences, increasing both memory and compute demands, as well as the end user’s cost by a factor of 2-3x when processing domain-specific text (Ahia et al., 2023). To mitigate this issue, prior work has explored expanding the token vocabulary during domain or language adaptation to improve tokenization efficiency (Wang et al., 2019; Zhao et al., 2024; Kim et al., 2024; Liu et al., 2023, 2024). While effective, this approach needs to be repeated for each target domain or language and requires maintaining separate tokenizers. Meanwhile, commercial LLM providers trend toward increasing the size of token vocabularies—growing from 32K to 128K (Grattafiori et al., 2024) and even up to 200K (Abdin et al., 2024) tokens—to improve overall tokenization efficiency. However, prior work (Dagan et al., 2024; Liang et al., 2023) shows that simply enlarging the vocabulary yields diminishing returns in domain adaptation, and vocabularies past a certain size can potentially degrade model performance (Liang et al., 2023).
These limitations point to a compelling need for an adaptive tokenization mechanism—one that can dynamically tailor the vocabulary to the input text at inference time, without retraining the model or maintaining separate tokenizers. Such a mechanism would allow the model to construct new domain-specific tokens on-the-fly, so to enhance tokenization efficiency. However, adaptive tokenization poses architectural challenges, as both the embedding layer and the language modeling head in transformer models (Vaswani et al., 2017) are static matrices tied to a fixed vocabulary size.
In this paper, we propose zip2zip (with a hat-tip to seq2seq (Sutskever et al., 2014)), a method that equips LLMs with a dynamic token vocabulary, enabling inference-time token adaptation. zip2zip achieves adaptive tokenization through continuous vocabulary expansion at runtime, allowing the model to represent a repeated or domain-specific pattern with a single long token rather than inefficient short tokens. This requires modest modifications to both the transformer architecture and the language modeling objective. zip2zip comprises three key components: (1) Tokenizer: an integration of Lempel-Ziv-Welch (LZW) compression [^1] LZW is the algorithm used in zip compression tool, which inspired the name zip2zip. (Welch, 1984) into the tokenization process, which continuously merges frequently co-occurring token sequences into reusable longer tokens (hypertokens) at runtime; (2) Architecture: a lightweight encoder added to the transformer that computes embeddings for newly formed tokens on the fly; (3) Training: a compression-aware causal language modeling variant that trains the model directly on LZW-compressed sequences, aligning learning with the inference-time token distribution. The name zip2zip reflects its dual role in achieving compression of both the input tokens (the first zip) and output tokens (the second zip), thereby jointly improving the efficiency of input encoding and output decoding. We finetune Phi-3-4B and Phi-3-14B to support zip2zip using as few as 100M tokens—requiring only 10 and 40 H100 GPU hours, respectively—for effective adaptation. The resulting models demonstrate strong inference-time compression capabilities and achieve 20–60% reductions in both input and output sequence lengths, translating to up to 60% improvements in end-to-end latency.
To make it easy to upgrade existing LLMs to zip2zip, we release an efficient, open-source implementation of the framework. It includes (1) a fast Rust-based LZW tokenizer, (2) a drop-in model architecture compatible with Hugging Face Transformers and vLLM, (3) a training pipeline for LZW-compression-based finetuning. Existing LLMs can be seamlessly extended with zip2zip, gaining adaptive tokenization capabilities through parameter-efficient finetuning—without any changes to the base model or tokenizer.
@@
Figure 1: Overview of the zip2zip inference pipeline. At each decoding step, the model has a growing context composed of both base tokens (blue) and hypertokens (green). The static vocabulary of size 6 remains fixed, while the dynamic vocabulary is continuously expanded by merging co-occurring tokens using LZW compression. The codebook (right) maps hypertoken IDs to their corresponding base tokens. As decoding progresses, new hypertokens created at step $t$ (e.g., “to be”, “or not”) become immediately available for reuse at step $t{+}$. Additionally, output tokens, once generated, instantly become eligible for compression. Hypertokens are also eligible for merging, enabling the formation of nested hypertokens. The final output sequence (bottom) is reconstructed via LZW decompression.
@@ 2.1 Dynamic Token Vocabulary
To enable dynamic tokenization at inference time, we associate each LLM with a hyper-vocabulary $\mathcal{V}_{h}$ that augments the model’s static token vocabulary. Tokens from the original vocabulary $\mathcal{V}$ are referred to as base tokens. Each entry in the hyper-vocabulary is a hypertoken, representing a merged sequence of base tokens. The total vocabulary for a zip2zip model is the union $\mathcal{V}\cup\mathcal{V}_{h}$. At the beginning of each inference session, $\mathcal{V}_{h}$ is initialized as an empty set, and is incrementally populated during decoding by identifying and merging recurring token subsequences in the context window, as illustrated in Figure 1.
@@
Continuous Vocabulary Expansion. As decoding proceeds, zip2zip continuously merge co-occurring tokens as new hypertokens to $\mathcal{V}_{h}$ and recurrently apply merging on newly generated tokens. This continual expansion allows the model to represent longer, recurring sequences of base tokens compactly. Hypertokens are treated as first-class tokens within the model, used interchangeably with base tokens throughout the decoding process. Importantly, this process occurs entirely during inference, without modifying the underlying tokenizer or requiring model retraining.
@@
LZW Algorithm. We implement vocabulary expansion using the Lempel-Ziv-Welch (LZW) compression algorithm—a dictionary-based, lossless compression method that incrementally builds a codebook of variable-length sequences. In our setting, the codebook is initialized with the base token vocabulary $\mathcal{V}$ and expands by adding new hypertokens on the fly as recurring token patterns are encountered. To control the growth of the dynamically expanding vocabulary, we impose a maximum merge size $M$ that restricts how many base tokens a single hypertoken can represent. LZW is particularly well-suited for zip2zip due to the following properties:
@@
     (1) it is streaming—hypertokens created at step $t$ can be immediately reusable at step $t+1$; in contrast, methods like BPE require access to the full sequence and operate offline;
@@
     (2) it is self-contained—input base tokens can be perfectly reconstructed from the compressed token sequence alone [^2] There is no need to persist or transmit the codebook across inference calls, preserving compatibility with existing LLM libraries and interfaces.;
@@
     (3) it is unambiguous—when both base tokens and hypertokens are available, which one to use is consistently determined by the LZW algorithm without ambiguity.
2.2 Hyper-Embedding and Hyper-Projection
@@
Hypertokens do not have fixed embedding vectors in the original model’s embedding layer (and projection layer), as they are not part of the original vocabulary. To compute the embedding of a hypertoken, we learn a mapping from the base token embeddings to the hypertoken embedding. We achieve this by introducing a hyper-encoder, which is a neural network that takes the embeddings of the constituent base tokens as input and outputs the corresponding hypertoken embedding. Specifically, for a sequence of $M$ base tokens $y_{1:M}:=y_{1}\dots y_{M}$, the hyper-encoder $f_{\phi}:\mathcal{V}^{M}\to\mathbb{R}^{d}$ produces the hypertoken embedding ${h}=f_{\phi}(y_{1:M})\in\mathbb{R}^{d}$, where $M$ is the maximum merge size and $d$ is the embedding dimension. For hypertokens composed of fewer than $M$ base tokens, we pad the input sequence to length $M$. Since the embedding map for base tokens remains unchanged, the hyper-encoder $f_{\phi}$ essentially maps the concatenated base token embeddings from a $(M\times d)$-dimensional space to a $d$-dimensional hypertoken embedding vector, performing nonlinear dimensionality reduction.
For the output projection layer, if the underlying transformer ties the embedding and the projection matrices, one can reuse the same hyper-encoder to compute the representation used for projection. Otherwise, a separate hyper-encoder is trained to produce the hypertoken projection vectors.
@@ 2.3 Architecture
We illustrate the architecture of zip2zip in Figure 2. The input text is first tokenized into base tokens (step 1), which are then passed through an online LZW compressing module that compresses the token sequence into a stream of hypertokens (step 2). Since hypertokens are not part of the model’s original embedding layer, their embedding vectors are computed on-the-fly using the hyper-encoder during inference (step 3–4). Once embedded, both base token embeddings and hypertokens embeddings are passed through the standard transformer layers of the base model, producing contextualized hidden states (step 5–6). This step is identical to vanilla transformer, with hypertokens and base tokens treated equally. At the output projection layer, hypertoken projection vectors (same as the hypertoken embedding vectors in the tied case, and computed by a separate hyper-encoder otherwise) are appended to the original projection matrix in the language modeling head (step 7). This allows the model to compute a joint softmax over the union of the base vocabulary and the hyper vocabulary $\mathcal{V}\cup\mathcal{V}_{h}$ (step 8). The resulting probability distribution is over $\mathcal{V}\cup\mathcal{V}_{h}$, and the sampled token may be either a base token or a hypertoken (step 9). In the next cycle, the newly generated token (step 10) —whether base or hyper—is appended to the input sequence, and the process repeats (back to step 1). At the end of generation, the hypertoken sequence is decompressed via the LZW decoding function into a sequence of base tokens (step 11–12).The whole process works in a fully autoregressive way, where newly generated tokens will also be merged into hypertokens for future steps. Furthermore, we highlight two points:
@@
Figure 2: zip2zip architecture. At inference time, base tokens are compressed into hypertokens using LZW (Steps 1–2). A hyper-encoder computes embeddings for hypertokens (Step 3–4), which are processed by the base LLM (Steps 5–6). Output representations are projected jointly on base and hyper-projection layers (Step 7), producing joint logits and sampled tokens (Steps 8–10), which can be decoded back to base tokens (Steps 11–12).
@@
Consistent Vocabulary Updates. The expanding vocabulary—comprising newly created hypertokens—must be updated in a consistent manner across both the input embedding layer and the output projection layer, maintaining a consistent view of the hypertoken set. Failure to update both sides consistently can result in two types of errors: (1) hypertokens that cannot be decoded, or (2) the model attempting to decode a non-existing hypertoken.
@@
Hyper-Embedding Cache. Although hypertoken embeddings are computed on-the-fly, they are context-independent and can thus be cached across inference steps. Similar to the transformer’s KV-cache, this enables incremental updates: only newly created hypertokens need to be processed at each step. Since the codebook grows linearly with the number of tokens in the context $n$, the total cache size grows also linearly in memory. Thus, the computational cost for hypertoken embeddings remains constant per step—i.e., one token embedding is computed per step.
2.4 Training zip2zip models
@@
Objective. Let $\mathcal{D}$ denote the target text distribution. Given a language model $\pi_{\theta}$ parameterized by $\theta$, standard pretraining seeks to minimize the causal language modeling (CLM) objective, which corresponds to the expected negative log-probability of data sequences under the model:
@@
To construct the compressed dataset $\mathcal{C}(\mathcal{D})$, we first tokenize the corpus using a standard tokenizer, and then apply the LZW compression algorithm. This preprocessing step is performed once prior to training and can be efficiently parallelized through batching.
@@
Auxiliary Reconstruction Loss. We introduce an auxiliary reconstruction objective that encourages a hypertoken embedding to retain sufficient information about its underlying base token sequence. Specifically, the model is trained to reconstruct the original base token embeddings from the hypertoken embedding. We jointly optimize the language model and the hyper-encoder using a combined loss that includes both the standard next-token prediction loss and the auxiliary reconstruction loss. Formally, we optimize:
@@
where $f_{\phi}:\mathcal{V}^{M}\to\mathbb{R}^{d}$ is the hyper-encoder, $f_{\psi}:\mathbb{R}^{d}\to\mathcal{V}^{M}$ is the decoder aiming to reconstruct the corresponding base tokens from their hyper-embedding, and $\Delta:\mathcal{V}^{M}\times\mathcal{V}^{M}\to\mathbb{R}$ is the reconstruction loss function, such as the cross-entropy loss, between the base tokens $y_{1:M}$ and the reconstructed base tokens $f_{\psi}\left(f_{\phi}(y_{1:M})\right)$. The hyperparameter $\lambda$ controls the trade-off between the prediction error of the language model and the reconstruction error of the autoencoder. This joint optimization objective encourages the hyper-encoder to learn a compact $d$-dimensional manifold embedded in the higher-dimensional $(M\times d)$ space of base token embeddings, while the language model $\pi_{\theta,\phi}$ learns to predict the next (hyper)token given the preceding context. The reconstruction loss can be viewed as a form of auto-encoding, where the hypertoken acts as a compressed latent representation and reconstruction encourages the preservation of semantic content and the compression to be lossless.
Adapting Pretrained Language Models. The proposed objectives (Equation 23) integrate naturally with pretrained language models. In this setting, the base model can be frozen while training only the hyper-encoder to adapt to compressed token sequences. Parameter-efficient methods such as LoRA (Hu et al., 2022) may also be used to adapt select components of the base model, enabling effective adaption with minimal computes.

### v2

# zip2zip: Inference-Time Adaptive Tokenization via Online Compression

@@ Abstract
Tokenization efficiency plays a critical role in the performance and cost of large language models (LLMs), yet most models rely on static tokenizers optimized on general-purpose corpora. These tokenizers’ fixed vocabularies often fail to adapt to domain- or language-specific inputs, leading to longer token sequences and higher computational costs. We introduce zip2zip, a novel method for achieving context-adaptive tokenization in LLMs at inference time. Leveraging an online data compression algorithm (Lempel–Ziv–Welch), zip2zip dynamically expands its active vocabulary at inference time by continuously replacing fragmented token sequences with more compact hypertokens, which it can immediately output during generation. zip2zip consists of three key components: (1) a tokenizer based on Lempel–Ziv–Welch compression that incrementally merges co-occurring tokens into reusable hypertokens on the fly; (2) a dynamic embedding (and unembedding) layer that computes embeddings for newly formed hypertokens at runtime; and (3) a variant of autoregressive language modeling that pretrains the model to handle hypertokenized, compressed text sequences as inputs and outputs. We show that an existing LLM can be uptrained for zip2zip in 10 GPU-hours via parameter-efficient finetuning. The resulting LLM performs test-time adaptation, learning to use hypertokens in unseen contexts and reducing input and output tokens by 15–40%. Code and models are released at https://github.com/epfl-dlab/zip2zip.
@@ 1 Introduction
Figure 1: zip2zip inference process. At each decoding step, the model has a growing context composed of both base tokens (blue) and hypertokens (green). The static vocabulary of size 6 remains fixed, while the dynamic vocabulary is continuously expanded by merging co-occurring tokens using LZW compression. The codebook (right) maps hypertoken IDs to their corresponding base tokens. As decoding progresses, new hypertokens created at step $t$ (e.g., “to be”, “or not”) become immediately available for reuse at step $t+1$. Hypertokens are also eligible for merging, enabling the formation of nested hypertokens. The final output sequence (bottom) is reconstructed via LZW decompression.
Large language models (LLMs) have shown impressive versatility across a broad spectrum of tasks and domains (Brown et al., 2020; Bubeck et al., 2023), including biomedical tests (Nori et al., 2023), mathematical reasoning (Frieder et al., 2023), programming (Jiang et al., 2024), and multiple human languages. A critical underlying component of this flexibility is the tokenizer, which defines the model’s vocabulary and governs how raw text is converted into token sequences fed to the model. The efficiency of the tokenization scheme—i.e., how compactly a text is represented as tokens—has significant impact on model performance. In particular, a more compact tokenization yields three key benefits: (1) larger effective context windows; (2) lower computational (and thus monetary) cost; and (3) shorter response times.
Despite its importance, the tokenizers used in most LLMs operate with fixed, static vocabularies obtained by running algorithms such as Byte Pair Encoding (Sennrich et al., 2016) over large-scale, general-purpose web corpora. While this globally optimized vocabulary performs reasonably well on average, it often fails to adapt to domain-specific or language-specific distributions (Ahia et al., 2023; Petrov et al., 2023), where the text distribution diverges significantly from the pretraining data. The resulting mismatch leads to longer token sequences, increasing both memory and compute demands, as well as the end user’s cost by a factor of 2–3x when processing domain-specific text (Ahia et al., 2023). To mitigate this issue, prior work has explored expanding the token vocabulary during domain or language adaptation to improve tokenization efficiency (Wang et al., 2019; Zhao et al., 2024; Kim et al., 2024; Liu et al., 2023, 2024a). While effective, this approach needs to be repeated for each target domain or language and requires maintaining separate tokenizers. Meanwhile, commercial LLM providers trend toward increasing the size of token vocabularies—growing from 32K to 128K (Grattafiori et al., 2024) and even up to 200K (Abdin et al., 2024) tokens—to improve overall tokenization efficiency. However, prior work (Dagan et al., 2024; Liang et al., 2023) shows that simply enlarging the vocabulary yields diminishing returns in domain adaptation, and vocabularies past a certain size can potentially degrade model performance (Liang et al., 2023). These limitations point to a compelling need for an adaptive tokenization mechanism—one that can dynamically tailor the vocabulary to the input text at inference time, without retraining the model or maintaining separate tokenizers. Such a mechanism would allow the model to construct new domain-specific tokens on-the-fly, so as to enhance tokenization efficiency. However, adaptive tokenization poses architectural challenges, as both the embedding layer and the language modeling head in transformer models (Vaswani et al., 2017) are static matrices tied to a fixed vocabulary size.
In this paper, we propose zip2zip (with a hat-tip to seq2seq (Sutskever et al., 2014)), a novel building block that brings inference-time adaptive tokenization to LLMs. zip2zip comprises three key components: (1) LZW tokenizer: A tokenizer that integrates the Lempel–Ziv–Welch [^1] LZW is the algorithm used in the ZIP compression tool, which inspired the name zip2zip. compression algorithm on top of Byte Pair Encoding (BPE) (Welch, 1984). By applying the LZW compression algorithm to the base token sequence—continuously merging frequently co-occurring token sequences into reusable longer tokens (hypertokens)—the resulting tokenization becomes less fragmented and more compact. (2) Dynamic-embedding architecture: An augmentation of the transformer architecture with a lightweight encoder that replaces the static embedding matrix, allowing the model to compute embeddings for newly formed hypertokens on the fly. (3) Pretraining under online token compression: a variant of causal language modeling that trains the model directly on LZW-compressed sequences, aligning learning with the inference-time (hyper)token distribution. The overall process is illustrated in Figure 1, which shows how the context window, dynamic vocabulary, and codebook evolve together during decoding. The name zip2zip reflects its dual role in achieving compression of both the input tokens (the first zip) and output tokens (the second zip), thereby jointly improving the efficiency of input encoding and output decoding. We conduct continued pretraining on Phi-3-4B and Phi-3.5-14B to support zip2zip using as few as 100M tokens. The resulting models demonstrate strong inference-time compression capabilities across various domains, achieving 15–40% reductions in sequence length and up to 40% improvements in end-to-end latency.
To make it easy to upgrade existing LLMs to zip2zip, we release an efficient, open-source implementation of the training and inference stack. It includes (1) a fast Rust-based LZW tokenizer, (2) a drop-in model architecture compatible with HuggingFace Transformers, (3) a training pipeline for LZW-compression-based finetuning. Existing LLMs can be seamlessly extended with zip2zip, gaining adaptive tokenization capabilities through parameter-efficient finetuning.
@@ 2.1 Dynamic Token Vocabulary
To enable dynamic tokenization at inference time, we associate the LLM with a hyper-vocabulary $\mathcal{V}_{h}$ that augments the model’s static token vocabulary. Tokens from the original vocabulary $\mathcal{V}$ are referred to as base tokens. Each entry in the hyper-vocabulary is a hypertoken, representing a merged sequence of base tokens. The total vocabulary for a zip2zip model is the union $\mathcal{V}\cup\mathcal{V}_{h}$. At the beginning of each inference session, $\mathcal{V}_{h}$ is initialized as an empty set, and is incrementally populated during decoding by identifying and merging recurring token subsequences in the context window, as illustrated in Figure 1.
@@
Continuous Vocabulary Expansion. As decoding proceeds, zip2zip continuously merges co-occurring tokens into new hypertokens and recursively applies merging on new hypertokens. This continual expansion allows the model to represent longer, recurring sequences of base tokens compactly. Hypertokens are treated as first-class tokens within the model, used interchangeably with base tokens throughout the decoding process. Importantly, this process occurs entirely during inference, without modifying the underlying tokenizer or requiring model retraining.
@@
LZW Algorithm. We implement vocabulary expansion using the Lempel–Ziv–Welch (LZW) compression algorithm—a dictionary-based, lossless compression method that incrementally builds a codebook of variable-length sequences. In our setting, the codebook is initialized with the base token vocabulary $\mathcal{V}$ and expands by adding new hypertokens on the fly as recurring token patterns are encountered. To control the growth of the dynamically expanding vocabulary, we impose a maximum merge size $M$ that restricts how many base tokens a single hypertoken can represent. LZW is particularly well-suited for zip2zip due to the following properties:
@@
  (1) it is online: hypertokens created at step $t$ can be immediately reusable at step $t+1$; in contrast, methods like BPE require access to the full sequence and operate offline;
@@
  (2) it is self-contained: input base tokens can be perfectly reconstructed from the compressed token sequence alone; [^2] There is no need to persist or transmit the codebook across inference calls, preserving compatibility with existing LLM libraries and interfaces.
@@
  (3) it is unambiguous: when both base tokens and hypertokens are available, which one to use is consistently determined by the LZW algorithm without ambiguity.
2.2 Hyper-Embedding and Hyper-Unembedding
@@
Hypertokens do not have fixed embedding vectors in the original model’s embedding layer (and unembedding layer), as they are not part of the original vocabulary. To compute the embedding of a hypertoken, we learn a mapping from the base token embeddings to the hypertoken embedding. We achieve this by introducing a hyper-encoder, which is a neural network that takes the embeddings of the constituent base tokens as input and outputs the corresponding hypertoken embedding (see Figure 2(a)). Specifically, for a sequence of $M$ base tokens $y_{1:M}:=y_{1}\dots y_{M}$, the hyper-encoder $f_{\phi}:\mathcal{V}^{M}\to\mathbb{R}^{d}$ produces the hypertoken embedding ${h}=f_{\phi}(y_{1:M})\in\mathbb{R}^{d}$, where $M$ is the maximum merge size and $d$ is the embedding dimension. For hypertokens composed of fewer than $M$ base tokens, we pad the input sequence to length $M$. Since the embedding map for base tokens remains unchanged, the hyper-encoder $f_{\phi}$ essentially maps the concatenated base token embeddings from an $(M\times d)$-dimensional space to a $d$-dimensional hypertoken embedding vector, performing nonlinear dimensionality reduction. For the output unembedding layer, if the underlying transformer ties the embedding and the unembedding matrices, one can reuse the same hyper-encoder to compute the representation used for unembedding. Otherwise, a separate hyper-encoder is trained to produce the hypertoken unembedding vectors.
Figure 2: (a) Dynamic embedding: Base tokens are embedded via a static LM embedding matrix, while hypertokens (e.g., “to be” or “to be or”) are dynamically composed using a hyper-encoder over their constituent base tokens. (b) Language modeling in compressed space: The model is trained to predict compressed token sequences produced by LZW, optimizing cross-entropy loss over compressed token IDs. (c) Auto-encoding loss: To ensure hypertokens are semantically consistent with their base-token compositions, the model also learns to reconstruct the original base tokens from the hyper-token via a decoding loss.
@@ 2.3 Architecture
Figure 3: zip2zip architecture and pipeline. At inference time, base tokens are compressed into hypertokens using LZW. A hyper-encoder computes embeddings for hypertokens, which are processed by the base LLM. Output representations are projected jointly on base and hyper-unembedding layers, producing joint logits and sampled tokens, which can be decoded back to base tokens.
We illustrate the zip2zip architecture in Figure 3. The input text is first tokenized into base tokens (step 1), which are then passed through an online LZW compressing module that compresses the token sequence into a stream of hypertokens (step 2). Since hypertokens are not part of the model’s original embedding layer, their embedding vectors are computed on-the-fly using the hyper-encoder during inference (step 3–4). Once embedded, both base token embeddings and hypertokens embeddings are passed through the standard transformer layers of the base model, producing contextualized hidden states (step 5–6). This step is identical to vanilla transformer, with hypertokens and base tokens treated equally. At the output unembedding layer, hypertoken unembedding vectors (same as the hypertoken embedding vectors in the tied case, and computed by a separate hyper-encoder otherwise) are appended to the original unembedding matrix in the language modeling head (step 7). This allows the model to compute a joint softmax over the union of the base vocabulary and the hyper vocabulary $\mathcal{V}\cup\mathcal{V}_{h}$ (step 8). The resulting probability distribution is over $\mathcal{V}\cup\mathcal{V}_{h}$, and the sampled token may be either a base token or a hypertoken (step 9). In the next cycle, the newly generated token (step 10)—whether base or hyper—is appended to the input sequence, and the process repeats (back to step 1). At the end of generation, the hypertoken sequence is decompressed via the LZW decoding function into a sequence of base tokens (step 11–12). The whole process works in a fully autoregressive way, where newly generated hypertokens will also be merged into new hypertokens for future steps. Furthermore, we highlight two points:
@@
Consistent Vocabulary Updates. The expanding vocabulary—comprising newly created hypertokens—must be updated in a consistent manner across both the input embedding layer and the output unembedding layer, maintaining a consistent view of the hypertoken set. Failure to update both sides consistently can result in two types of errors: (1) hypertokens that cannot be decoded, or (2) the model attempting to decode a non-existing hypertoken.
@@
Hyper-Embedding Cache. Although hypertoken embeddings are computed on-the-fly, they are context-independent and can thus be cached across inference steps. Similar to the transformer’s KV-cache, this enables incremental updates: only newly created hypertokens need to be embedded at each step. Since the codebook grows linearly with the number of tokens in the context, the total cache size also grows linearly in memory. Thus, the computational cost for hypertoken embeddings remains constant per step—i.e., one token embedding is computed per step.
2.4 zip2zip Pretraining
@@
Objective. Let $\mathcal{D}$ denote the target text distribution. Given a language model $\pi_{\theta}$ parameterized by $\theta$, standard pretraining seeks to minimize the causal language modeling (CLM) objective (see Figure 2(b)), which corresponds to the expected negative log-probability of data sequences under the model:
@@
To construct the compressed dataset $\mathcal{C}(\mathcal{D})$, we first tokenize the corpus using a standard tokenizer, and then apply the LZW compression algorithm. This preprocessing step is performed once prior to training and can be efficiently parallelized through batching. Compression is applied at the document level, meaning that each document is compressed independently. This prevents the compressor from learning patterns across unrelated documents.
@@
Auxiliary Reconstruction Loss. We introduce an auxiliary reconstruction objective that encourages a hypertoken embedding to retain sufficient information about its underlying base token sequence (see Figure 2(c)). Specifically, the model is trained to reconstruct the original base token embeddings from the hypertoken embedding. We jointly optimize the language model and the hyper-encoder using a combined loss that includes both the standard next-token prediction loss and the auxiliary reconstruction loss. Formally, we optimize:
@@
where $f_{\phi}:\mathcal{V}^{M}\to\mathbb{R}^{d}$ is the hyper-encoder, $f_{\psi}:\mathbb{R}^{d}\to\mathcal{V}^{M}$ is the decoder aiming to reconstruct the corresponding base tokens from their hyper-embedding, and $\Delta:\mathcal{V}^{M}\times\mathcal{V}^{M}\to\mathbb{R}$ is the reconstruction loss function, such as the cross-entropy loss, between the base tokens $y_{1:M}$ and the reconstructed base tokens $f_{\psi}\left(f_{\phi}(y_{1:M})\right)$. The hyperparameter $\lambda\geq 0$ controls the trade-off between the prediction error of the language model and the reconstruction error of the autoencoder. This joint optimization objective encourages the hyper-encoder to learn a compact $d$-dimensional manifold embedded in the higher-dimensional $(M\times d)$ space of base token embeddings, while the language model $\pi_{\theta,\phi}$ learns to predict the next (hyper)token given the preceding context. The reconstruction loss can be viewed as a form of auto-encoding, where the hypertoken acts as a compressed latent representation and reconstruction encourages the preservation of semantic content and the compression to be lossless.
Adapting Pretrained Language Models. Retraining large language models from scratch is computationally expensive and often infeasible for most research labs. A more economical alternative is to perform continued pretraining (or adaptation) on existing pretrained model weights. The proposed objectives (Equations 23) integrate naturally into this setup. Parameter-efficient methods such as LoRA (Hu et al., 2022) may also be used, which allow selectively updating parts of the base model weights with minimal computational cost.